# Lightning declines over shipping lanes following regulation of fuel sulfur emissions

### v3

@@ Acknowledgements
We use the GPM precipitation feature database from Chuntao Liu’s group to probe the shipping lane enhancement of cold, strong reflectivity echoes (doi.org/10.1175/2008JAMC1890.1). Note that this approach to examining reflectively is slightly different from that done in Thornton et al., (2017), but follows that in Blossey et al., (2018). Due to infrequent sampling and small swath width, the reflectivity signal shown here requires very long records to show the effect of the shipping lane (previous studies, such as Blossey et al., (2018), use the full TRMM record, which is well over a decade). Splitting the relatively short (8 year) GPM record of reflectivity into a pre- and post-IMO period therefore greatly reduces the statistical power of the retrievals. Indeed, the peak in reflectivity established over the shipping lane between 2016 and 2019 has diminished (Figure S5). However, the difference between the two periods is not statistically significant (p less than 0.05), and we therefore refrain from concluding that there has or has not been a change to reflectivity above the shipping lanes.

### v4

@@ Acknowledgements
Thornton et al (2017) showed that DCC over the shipping lane exhibited anomalously higher radar reflectivity in the mixed phase region prior to 2015, but we find statistically robust tests of changes in DCC reflectivity since the IMO regulation are not yet possible. We use the GPM precipitation feature database from Chuntao Liu’s group to probe the shipping lane enhancement of cold, strong reflectivity echoes (doi.org/10.1175/2008JAMC1890.1). Note that this approach to examining reflectively is slightly different from that done in Thornton et al., (2017), but follows that in Blossey et al., (2018). Due to infrequent sampling and small swath width, the reflectivity signal shown here requires very long records to show the effect of the shipping lane (previous studies, such as Blossey et al., (2018), use the full TRMM record, which is well over a decade). Splitting the relatively short (8 year) GPM record of reflectivity into a pre- and post-IMO period therefore greatly reduces the statistical power of the retrievals. Indeed, the peak in reflectivity established over the shipping lane between 2016 and 2019 has diminished (Figure S5). However, the difference between the two periods is not statistically significant (p less than 0.05), and we therefore refrain from concluding that there has or has not been a change to reflectivity above the shipping lanes.